### v2

@@ Abstract
For example, one may want to guide an inverse-folding model to produce proteins with greater solubility than the model naturally generates, or to produce proteins more likely to bind a particular target, and so forth. Perhaps more importantly, one might want to guide a model using newly acquired experimental measurements, such as catalytic activity for a desired reaction—that is, integrate experimental data in a statistically coherent manner with an existing generative model, without re-training the existing model. Such an ability would enable training-efficient iterative cycles of design, build, test, and learn—a process where experimental feedback guides subsequent AI-based design rounds.
We recently showed for the first time how to correctly and tractably perform guidance for two specific classes of generative models, diffusion and flow matching, on discrete state-spaces [16]. Discrete state-spaces include not only sequences of amino acid, DNA and RNA, but also SMILES strings, graphs, language tokens, structure tokens, or any object represented by a sequence of discrete tokens. Herein, we build on this work by i) generalizing our guidance method to two more commonly-used classes of generative models for discrete spaces, namely, a large class of Masked Language Models (MLMs) such as ESM3, and order-agnostic autoregressive models (OA-AR) such as ProteinMPNN, yielding our new method, ProteinGuide, ii) illustrating in silico application of ProteinGuide to representative use cases in protein engineering, and iii) applying ProteinGuide to obtain adenine base editors showing enhanced activity experimentally compared to the unguided model.
Figure 2: In silico and experimental demonstration of ProteinGuide. a. Generating structure-conditioned protein sequences from ProteinMPNN without guidance (teal) and with ProteinGuide for stability (pink), across eight proteins (horizontal axis). The height of each bar represents the percentage of sequences predicted to be at least as stable as wild type ($\Delta\Delta G\leq 0$), while the hatched sub-portion indicates the fraction of those stable sequences that are also predicted to correctly fold into the desired structure (RMSD $\leq 2$Å). Error bars on all panels represent standard errors of proportions, calculated as $\sqrt{p(1-p)/n}$, where $p$ is the empirical proportion of success and $n$ is the sample size. b. Using ESM3 to re-design enzyme sequences without guidance (teal) and with ProteinGuide (pink). The % residues re-designed refers to how many residues (positions) of a randomly chosen enzyme sequence from the target class were masked for re-design. When all are masked (100%), the task becomes de novo design. Bar height indicates the percentage of generated samples classified by CLEAN [17] as being from the correct (specified) target enzyme class. The hatched sub-portion shows the fraction of correctly classified samples with median pLDDT $>0.8$. c. Generating backbone structure tokens with ESM3 without guidance (teal), and with fold class ProteinGuided ESM3 (pink), across different CATH hierarchy levels. The height of each bar shows the percentage of structures correctly classified to their target fold class, while the hatched sub-portion indicates the subset that are also designable (self-consistency RMSD $\leq 2$Å). Results are grouped by CATH level. For level C, there are three classes: mainly $\alpha$, mainly $\beta$, and mixed $\alpha/\beta$. For level A and T, we grouped the classes for each level by their frequency of occurrence in the AFDB (Common, Regular, Rare) and calculated the metrics within each group following Geffner et al. [18]. d. Schematic representation of the ProteinGuide ABE experiment. In round 1 we diversified a 86-amino acid TadA region using two generative models, ESM3 and FMIF, and built a library of 2,000 variants (library 1) which was assayed in a bacterial selection strain. The estimated activities of the library members were used to train a classifier to ProteinGuide for a second round of designs. This second round, library 2, contained 2,000 sequences that were cloned and assayed (round 2) and the activities of the two libraries were compared. e. A histogram of the experimentally measured log enrichments for both library 1 (teal) and library 2 (pink). ProteinGuide was used to generate the round 2 library such that variants would be enriched for having high activity (predicted activity in log enrichment units larger than the gray dashed line).
The essential technical formalism required to derive ProteinGuide, namely to generalize guidance in Nisonoff et al. [16] to MLMs and OA-AR models, is the connection between how these models and diffusion models are trained. Although conceptually diffusion/flow matching models, MLMs, and OA-AR models seem to be quite different, it turns out that the loss functions used for training them are equivalent (Fig. 1a and Methods). [^2] For MLMs, the equivalence holds when the masking rate at training time covers fully masked, fully unmasked, and everything in between, such as for ESM3 (see Methods). Consequently, we can view a trained MLM or OA-AR model as a diffusion model on a discrete state-space, allowing us to leverage our earlier guidance method [16] for these models (Fig. 1b and Methods). While connections between these training objectives has been previously noted [19, 20, 21, 22, 23], the practical utility of this observation has been unclear. Herein, we show that this connection enables a general framework for guiding a broad class of models. Note that for autoregressive (AR) models, one can use FUDGE [24] for guidance, which can be viewed as a special case of ProteinGuide for the case of AR models. More detailed discussion of related work can be found in the Methods.
[…]

### v4

@@ Abstract
ProteinGuide can be used, for example, to guide an inverse-folding model to produce proteins with greater solubility/expression/activity than the model naturally generates, or, to produce proteins more likely to bind a particular target, and so forth. Predictive models for generic properties can be used, such as those for stability/expression/solubility, or new, user-created models trained on newly acquired experimental measurements, such as catalytic activity for a desired reaction. Owing to the latter, ProteinGuide enables easy integration of experimental data in a statistically coherent manner with an existing generative model, without re-training the existing model. Consequently, ProteinGuide should be a key enabler in training-efficient iterative cycles of design, build, test, and learn—a process where experimental feedback guides subsequent AI-based design rounds.
Part of the technical underpinnings for ProteinGuide rely on our earlier development of guidance for continuous-time diffusion and flow-matching models on discrete state-spaces, referred to herein as Discrete Guidance [82]. Following on that work, others found that Discrete Guidance performed favorably compared to fine-tuning for protein property design tasks when used to guide their family-specific generative models [125]. Crucially, Discrete Guidance did not address guidance for models other than discrete diffusion models (DDMs) [4, 12, 74] and discrete flow matching models (DFMs) [13, 32], thereby limiting its utility to the field of protein engineering and beyond. For example, our earlier work did not provide a way to guide ESM3 [44], ProteinMPNN [20] and many other models [76, 81, 31]. With ProteinGuide, we extend guidance to also include masked language models with masking rates covering the range $0\%-100\%$, sometimes known as generative masked language models (MLMs) [36, 115, 14, 68], and any-order autoregressive models (AO-ARMs) [112, 127, 49]. In so doing, ProteinGuide unlocks guidance to much larger set of models that are actively being developed by the protein engineering community.
To demonstrate the potential utility of ProteinGuide, we deployed it on a range of problems in protein engineering, first evaluating the approach in silico, and then in the wet lab. Specifically, we demonstrate the applicability of ProteinGuide in silico by guiding different pre-trained protein generative models, including ProteinMPNN, ESM3, and ESM C [31], to generate amino acid and structure token sequences, conditioned on several user-specified properties. These properties include higher stability and activity, functional annotations such as enzyme classes and CATH-labeled folds. We show results for generating new sequences with similar properties to the training data (interpolative), and also pushing beyond that (extrapolative). We also demonstrate that ProteinGuide can be used on more challenging problems wherein we seek to design for two different desired properties that are in tension with one another (increased on-target, and decreased off-target activity). Finally, we demonstrate the utility of our method experimentally by using ProteinGuide with two different inverse folding models to design new adenine base editors. We obtained experimental editing activity measurements in a high-throughput assay and train a predictive model to design sequences with higher editing activities. We show that editing efficiency previously achieved with seven rounds of directed evolution can now be surpassed with just a single round of guided design with ProteinGuide.
Results
We begin by introducing the concept of guidance and the theoretical framework underlying ProteinGuide, from which we derive the guidance algorithms used throughout this work. We then place ProteinGuide in the context of alternative strategies for incorporating auxiliary information into generative models, before presenting our in silico and in vivo experimental results.
Figure 1: Overview of ProteinGuide. a) Contrasting unguided generation with guided generation by way of ProteinGuide. A sequence is denoted, $x$, the pre-trained generative model is denoted $p_{\theta}(x)$; $y$ denotes a property of interest that we seek to guide toward using predictive model $p_{\phi}(y\mid x)$. When guiding, we blend information from the predictive model with that of the pre-trained model in a statistically correct manner, resulting in sequences sampled from the property-informed, reweighted distribution, $p_{\theta,\phi}(x\mid y)$. Masking of one sequence position is denoted with the symbol “?”, and pink letters indicate key mutations that are introduced only when guiding with the predictive model, rather than when using unguided generation. b) Unifying view of training and sampling for masked diffusion/flow models, MLMs, and AO-ARMs enables guidance on a broad set of model classes. Colored dots visually tag individual sequences, each of which has a different amount of masked tokens. The sequence being masked during training is M K T R S V. The shared training objective involves randomly masking out parts of the sequence and then predicting the masked sequence given the remaining unmasked context over different masking rates. Once any type of model is trained, one can then use a sampling algorithm originally developed for any of the other models. Consequently, we can obtain computationally tractable exact sampling for continuous time models and also accelerated approximate sampling. c) Schematic contrasting ProteinGuide with post hoc filtering and fine-tuning. ProteinGuide combines a pre-trained generative prior with assay-labeled data through correct statistical conditioning, enabling more reliable enrichment for high-property sequences than the other approaches shown.
Intuitive overview of guidance
We here begin our intuitive overview of ProteinGuide with the notion of guidance itself. Conceptually, performing guidance comprises taking i) a pre-trained generative model, $p(x)$, on sequences, $x$, and ii) a classifier/regression predictive model, $p(y|x)$ for property $y$, to enable sampling (i.e., generation) from the desired conditional generative model, $x\sim p(x|y)\propto p(y|x)p(x)$. Notably, guidance does so without training/re-training an underlying generative model on xx—rather, it modulates the sampling process of the unconditional model. Specifically, we execute Bayes’ Rule at each sampling step to compute a property-conditioned distribution for the next amino acid using the unconditional model and the predictive model. This makes the guided sequence an effective sample from a property-conditioned generative model without ever needing to materialize the conditional generative model itself. We therefore refer to guidance as “on-the-fly” conditioning. It does not require training a generative model—it only modulates the sampling process of the unconditional model.
We first develop guidance for discrete-space diffusion and flow-matching models, a method we earlier dubbed Discrete Guidance [82], which builds on prior work for continuous-space diffusion models. In continuous spaces, the key insight enabling guidance was that the generative process uses a series of gradient steps, each of which can be tractably converted to a conditional gradient with Bayes’ Rule. Consequently, the corresponding sequence of gradient steps enables sampling from the desired conditional distribution [100, 101, 26]. However, for discrete-space generative models, such as discrete-space diffusion and flow-matching models, these gradients are not defined. In fact, in discrete spaces, such models need to be set up completely differently even without guidance, to rely on transition rates instead of gradients [12, 107, 74, 13, 32]. In an analogous manner to deriving conditional gradients for continuous-space diffusion guidance, we derived conditional transition rates to unlock guidance for diffusion and flow-matching models on discrete spaces such as sequences (Section 1.2).
Our derivation shows that to sample from the desired conditional generative model, one modulates the unguided transition rates with a likelihood ratio of the desired property condition before and after each possible transition (Section 1.2). The key to enabling practical application of guidance with these conditional rates is to leverage aspects of the continuous-time process to enable what is naively an intractably expensive computation, to one that is manageable. Specifically, we show how to change naive guidance from being combinatorial in the sequence length at each sampling step, to just linear, by noticing that it is sufficient to assume that only one position mutates in infinitesimal time (Section 1.2). We also show how to speed up this exact guidance even further by approximating repeated calls to the predictive model with a first order approximation to the predictive model. This even faster, approximate guidance, called Taylor-approximate guidance (TAG), often matches the sample quality of exact guidance on many tasks [82].
Overview of ProteinGuide and theoretical results
ProteinGuide is a unifying framework that extends Discrete Guidance to a broad range of discrete-space generative models by exploiting equivalences in their training objectives and sampling distributions shown herein (Section 1.3 and 1.4). A key observation is that when diffusion and flow-based models employ a so-called masking noise process—a common choice for protein sequence modeling—their training objectives and sampling distributions reduce to forms that are mathematically equivalent to those of generative MLMs and AO-ARMs (Section 1.3 and 1.4). As a result, four widely used model classes—DDMs, DFMs, MLMs, and AO-ARMs—can be unified under a single framework comprising two regimes: continuous-time models (DDMs and DFMs) and discrete-time models (MLMs and AO-ARMs) (Fig. 1b). By establishing their training and sampling equivalences, ProteinGuide allows guidance algorithms to be transferred across model classes, enabling any of the four to be guided using either a continuous-time (DFM-style) or discrete-time (AO-ARM-style) formulation.
Although these two strategies are theoretically equivalent, they have practical differences that give rise to two different concrete algorithms. The first is Taylor-approximate guidance (TAG), a fast, approximate method proposed in Discrete Guidance that we use to guide pre-trained generative models as DFMs using the derived training equivalences. We used TAG in the majority of our experiments. The second is the newly derived method that we called discrete-time exact guidance (DEG). This is a slower, but exact, guidance algorithm that uses our sampling equivalence to directly extend Discrete Guidance to discrete-time models. We use DEG to guide models as AO-ARMs when either the numerical integration used to sample from a continuous time model is unstable or the first-order approximation used by TAG is insufficient. In particular, we used DEG only in our multi-property Pareto-extrapolative and enzyme class guidance experiments.
While connections between the training objectives of aforementioned model classes have been previously noted in various forms [4, 49, 98, 94, 83], the practical utility of these observations has been unclear. Herein, we show that this connection can be used to guide MLMs and AO-ARMs by treating them as discrete flow matching models so as to then apply Discrete Guidance. We also note that our sampling equivalence can be used more generally to define equivalent AO-ARMs to exactly sample from any DFM or DDM, with or without guidance. This generalizes recent results for DDMs showing that efficient “Gillespie” samplers could be created for those models [86, 17, 49, 132, 83, 98, 94], and shows, for the first time, that exact sampling from a DFM can be done tractably. Previously this was assumed to be intractable because it would require infinitesimally small numerical integration time steps [13, 82, 43].
Altogether, our unifying view of training and sampling brings these four model classes under a single framework and enables ProteinGuide to guide a broad set of discrete-space generative models, making it readily applicable to many popular protein sequence generative models commonly used by practitioners (Fig. 1b).
Alternative approaches for infusing generative models with auxiliary information
The simplest approach for generating sequences from pre-trained models with desired property criteria is to post hoc filter the generated sequences to only those satisfying the user-specified criteria (Fig. 1c). Such an approach can be viable when the pre-trained model by itself frequently generates sequences satisfying the criteria. However, more generally, such an approach is extremely inefficient because the properties of interest are not well-captured by general pre-trained models. We shall also see this in our experiments. Any approach that stands a chance of outperforming post hoc filtering must involve, in some fashion, “adapting” the generative models to infuse it with knowledge of the relationship between sequences and the desired properties, as ProteinGuide does.
Perhaps the most commonly used approach in protein engineering for infusing experimental data into a generative model is by way of so-called “fine-tuning” (FT, sometime referred to as supervised finetuning, see Section 1.5[133, 84]. FT involves training the generative model further with its original objective on a smaller, curated set of examples that have the desired properties (Fig. 1c). In the context of protein sequence generative models, this set might comprise sequences from a particular family that one hopes to focus the model on [92, 81, 39, 80, 15], or sequences that are categorized as being functional [8, 104]. FT has the benefit of being immediately applicable to any class of generative models. However, it has two downsides. First, FT is a heuristic method for combining sources of information, for which the statistical outcome is not well-defined (as opposed to say ProteinGuide which executes Bayes’ Rule). Consequently, important information from the pre-trained model may be lost, a phenomenon often referred to as “forgetting” [64, 75]. Second, the act of continuing to train the original objective on new data concentrates the existing pre-trained distribution onto the portion corresponding to the new data, limiting generation to sequences whose properties resemble those seen during training [81, 80, 92]. We see evidence of this in its relative inability to extrapolate in several of our in silico experiments (see Section “Extrapolative experiments: guiding ESM3 to generate sequences from a family for desired properties” and “Multi-property, Pareto-extrapolative experiment: guiding ESM C for on- and off-target binding of PbrR”). Nevertheless, FT has been shown to be useful either on its own, or as a first step before applying other approaches [133, 84, 8] and is likely sufficient for designing proteins with similar property values to the training data.
Fundamentally, guidance, and consequently ProteinGuide, differs from both post hoc filtering and fine-tuning in that it uses a stand-alone predictive model to infuse auxiliary information into the pre-trained model, and does so using the statistical operation of executing Bayes Rule. Consequently, ProteinGuide provides a framework with which to understand how we are blending sources of information. Additionally, because the predictive model can have class probabilities, or real-valued predictions, it can provide more nuanced information than FT which is limited to using only a select set of samples without associated property strengths. Because ProteinGuide executes Bayes’ rule, it can coherently blend several independent predictive models to simultaneously condition on all properties (see Section “Multi-property, Pareto-extrapolative experiment: guiding ESM C for on- and off-target binding of PbrR”).
Other statistically grounded frameworks have also been proposed. For example, Direct Preference Optimization (DPO) [87] and related work [120, 18, 117, 89] enable Bayesian-style updating of a pre-trained model using “preference-labeled” data or user specified “reward function”. Notably, these methods all require training a new generative model before one can sample from it. Up until recently, DPO has been applicable only to models with tractable likelihoods, thereby restricting them largely to autoregressive models [120], although DPO was recently extended to diffusion models on continuous spaces by way of approximation [114], and then to discrete spaces [9]. Further research is needed to understand and improve the quality of such approximations. Herein, we focus our study on “on-the-fly” guidance, rather than approaches that require training a generative model. However, we believe that such approaches will also prove useful.
Our view is that guidance offers an approach to statistical conditioning that is theoretically sound, practically well-behaved, flexible, and complementary to existing techniques. Importantly, guidance uniquely does not require access to the parameters of pre-trained models—it only requires predictions (logits) from the model. Accordingly, it can be used with closed-source models and large open source models for which fine-tuning might not be feasible for many users. Additionally, in settings where useful property-predictive models already exist, guidance allows one to directly leverage those models for the design task at hand. In recent work comparing guidance-based approach to fine-tuning for protein property-focused design tasks, it was found that our Discrete Guidance [82] performed favorably compared to fine-tuning based approaches, and that guidance requires less hyperparameter tuning and computation, making it more practically accessible to everyday users than other approaches examined [125]. For further discussion of related work, including that of Markov Chain Monte Carlo (MCMC) methods [30, 122, 67, 99], we refer the readers to Section 2.1.
Overview of in silico experiments
To illustrate the potential of ProteinGuide, we first applied it, in silico, to a suite of experiments spanning three design settings, using a set of representative, well-known protein generative models, including ProteinMPNN [20], ESM3 [44], and ESM C [31], as illustrative examples. See Supplementary Table S1 for an overview of these experiments. Note that ProteinGuide can also be applied to many other sequence generative models [1, 118, 13, 76, 81].
The three design settings capture three valuable use-cases. The first is an interpolative design setting, wherein we sought to generate proteins with property values that were seen in the supervised training data, but with greater diversity than appeared there. In a second category, an extrapolative design setting, we sought to generate diverse proteins with property values that exceeded any seen in the supervised training data. Finally, in a third task category, multi-property, Pareto-extrapolative, we sought to improve one property, while decreasing another, such that we could push beyond the optimality frontier of these two properties defined by the training data for the supervised model.
For the interpolative task, we guided the inverse-folding model ProteinMPNN with a stability predictive model to generate sequences. We also guided the multi-modal model ESM3 with function annotations to generate sequences and structures. Because the interpolative setting is a relatively easy one for design, we highlight the former in the main text, leaving the other to the Supplementary Information (Section 1.91.10, Supplementary Figure S6S7).
For the extrapolative tasks, we guided ESM3 to generate sequences for each of three proteins, with predictive models trained on assay-labeled data measuring fluorescence, ubiquitin ligase activity, or in vivo functional complementation. For the extrapolative, multi-property Pareto task, we guided ESM C to generate PbrR sequences, that simultaneously bind the desired lead ion with higher affinity and the off-target, zinc ion with low affinity.
The interpolative tasks had hundreds of thousands of data points with which to train the predictive model, while the extrapolative tasks had around 2,000, and the Pareto-extrapolative task had around 1,000. These numbers were chosen to be in line with our wet lab experiment where we used 2,000 assayed variants to train our predictive model.
In all of these in silico experiments, we compared ProteinGuide to the most commonly-used approaches in practice, namely, fine-tuning and post hoc filtering. We used LoRA [54] when performing fine-tuning with ESM3 and ESM C. In addition to these two baselines, we compared to the pre-trained model without any guidance or other changes. This baseline demonstrates how frequently the desired property would emerge by simple use of a general, unconditional model, reflecting the difficulty of the task for current models. For meaningful comparison, we allotted each method the same fixed compute budget as measured by wall-clock time, and a sample budget of $k$ samples. Full experimental details, including datasets, predictive models, baseline configurations, sampling settings, as well as additional results, are provided in the Methods and Supplementary Information. We now go through the in silico experimental results.
Interpolative experiments: guiding ProteinMPNN for improved stability
Here we illustrate how assay-labeled experimental data can be used to condition protein sequence generative models to improve their success in generating sequences with a desired function, as measured in those data. We used ProteinGuide to guide ProteinMPNN, an AO-ARM, with a predictive model trained on a large dataset of experimental folding stability measurements of mini-proteins [108]. The goal was to generate sequences with greater stability than the wild-type (Fig. 2a). A similar experiment was previously conducted by re-training the inverse folding model, which ProteinGuide lets us bypass [120].
We conducted these experiments on eight randomly chosen proteins from the stability dataset for which sequences generated by ProteinMPNN were less stable than the wild-type sequence on average. For each protein, the success rate was computed over $k=100$ generated sequences, where success was defined based on two criteria: (i) the generated sequence should fold into the desired structure, and (ii) the generated sequence should be at least as stable as the wild-type sequence, $\Delta\Delta G\leq 0$. We used ESM3 to predict the structure of the generated sequences and compare the root mean square deviation (RMSD) between the predicted structure and the desired structure; we considered the predicted designed structure to be successful if the RMSD was less than or equal to 22Å, following previous evaluation criteria [13]. To evaluate the folding stability of the generated sequences, we used the stability oracle from Nisonoff et al. [82].
In addition to ProteinMPNN’s default autoregressive sampling, through our equivalences, we could also more quickly sample from it as a flow-matching (FM) model, which we also compared to. In implementing the FT baseline, we followed a similar procedure to Widatalla et al. [120] in that we further trained ProteinMPNN on its original objective on all sequences more stable than its corresponding wild-type in the stability dataset (Section 1.6).
Across all eight proteins, ProteinGuide consistently yielded the highest success rates (Fig. 2b). In particular, the generated sequences from ProteinGuide preserved the structure constraints of the backbone as measured by refolded RMSD (Fig. 2c), while simultaneously improving their predicted stability (Fig. 2d). As expected, post hoc filtering worked best when the unguided model had at least a moderate success rate. FT improved success rates of the unguided model in some cases, but fails to do so in others, for reasons that are not yet fully understood. In contrast to these two methods, ProteinGuide worked well across the board. Note that diversity and novelty of generated sequences was similar across all methods (Supplementary Figure S1).
Figure 2: Guiding ProteinMPNN with experimental stability data to generate sequences with enhanced stability. a) Schematics for guiding ProteinMPNN with large-scale experimental stability measurements. ProteinMPNN, an inverse folding model, is guided by a stability predictive model trained on the large-scale stability dataset from Tsuboyama et al. [108] to generate sequences with stability better than the wild-type sequence. b) Five methods for sampling (in legends) are assessed across eight proteins (horizontal axis) by their success rate on 100 generated samples. The height of each bar represents the percentage of sequences predicted to be at least as stable as wild type ($\Delta\Delta G\leq 0$), while the hatched sub-portion indicates the fraction of those stable sequences that are also predicted to correctly fold into the desired structure (RMSD $\leq 2$Å). In most cases the hatched sub-portion covers the full height of the bar. Error bars represent standard errors of proportions, calculated as $\sqrt{p(1-p)/n}$, where $p$ is the empirical proportion of success and $n$ is the sample size. c) Distributions of refolded RMSD for the generated sequences (lower is better). The dashed horizontal line indicates the threshold below which a sequence is considered to fold into the desired structure (RMSD $\leq 2$Å). d) Distributions of oracle predicted stability for the generated sequences (lower is better). The dashed horizontal line indicates the threshold below which a sequence is predicted to be at least as stable as the wild-type ($\Delta\Delta G\leq 0$).
Extrapolative experiments: guiding ESM3 to generate sequences from a family for desired properties
Having investigated the interpolative design setting, we next evaluated ProteinGuide with the more challenging task of generating sequences with higher property values than those measured in the assay-labeled datasets (Fig. 3a). We used assay-labeled datasets for three different proteins, each one having had a distinct property measured, and each data set including many variants beyond single mutations (Supplementary Figure S2a). Specifically, we used data from i) the CreiLOV fluorescent protein, where the assay measures fluorescence intensity [16], ii) the ubiquitination factor E4B (UBE4B), where the assay measures E3 ubiquitin ligase activity [102], and iii) the poly(A)- binding protein Pab1 (PABP), where fitness is measured by way of an in vivo functional complementation selection [78].
For our pre-trained model, we used ESM3 unconditionally. For each of the three proteins, we guided generation with a predictive model trained on 2,000 randomly chosen labeled sequences from that protein family. Following similar experimental set-up to prior work [8, 125, 30], we used an “oracle” model trained on all assay-labeled dataset as an in silico proxy to assess the fitness of the generated sequences (Supplementary Figure S2b). Since ESM3 is a “pan-protein” model, we also simultaneously guided ESM3 to generate sequences belonging to the desired protein family. This required training a predictive model to discriminate between sequences in the assay-labeled dataset and unguided generations from ESM3 (Section 1.7).
For each protein design task, we allowed a budget of $k=100$ sequences. For fine-tuning in this set of experiments, we trained on the $M$ top variants from the assay-labeled data set. The value of $M$ was chosen by scanning through different $\%$ threshold on top-scoring variants as a hyperparameter, and choosing the one that had highest likelihood for held-out high-scoring variants.
Across all three proteins, ProteinGuide consistently yielded the highest success rates. In particular, ProteinGuide generated sequences with highest oracle property values compared to other methods (Fig. 3b), and with higher diversity and novelty than those generated with FT (Fig. 3c). Moreover, some of ProteinGuide’s variants had higher property values than any observed in the assay-labeled dataset (Fig. 3d). As expected, FT samples’ oracle values recapitulated the range of those in its training dataset (Fig. 3c,d). We hypothesize that this occured because FT concentrates the distribution of the pre-trained model on the new training data—useful for interpolative design. ProteinGuide, on the other hand, has the potential to extrapolate beyond the observed dataset to the extent that the predictive model used for guidance is able to extrapolate.
Figure 3: ProteinGuide guides ESM3 with assay-informed predictive models for extrapolative design. a) Experimental set-up for the extrapolative setting, where the goal is to generate sequences with higher fitness compared to those in the labeled dataset. The set of sequences (around 100) with highest property values in the labeled dataset were excluded for both fine-tuning and predictive model training. Fine-tuning is performed on sequences on $M$ top variants from the assay-labeled data set. $M$ is chosen for each dataset scanning through different % threshold on top-scoring variants as a hyperparameter, and choosing the one that had highest likelihood for held-out high-scoring variants. The predictive model used in guidance is trained on 2,000 sequences randomly sampled from the labeled dataset. b) For each dataset (subpanel title), each method is scored by its success rate on 100 samples. The height of each bar represents the percentage of sequences predicted by the oracle model to be above a particular threshold based on the quantile of the training set (horizontal axis) and are not identical to any sequence in the labeled set (i.e., novelty >0>0), while the hatched sub-portion indicates the fraction of those sequences that are also predicted by ESMFold [69] to fold into a structure with pLDDT $\geq$ 0.8 (Supplementary Figure S3). Error bars represent standard errors of proportions, calculated as $\sqrt{p(1-p)/n}$, where $p$ is the empirical proportion of success and $n$ is the sample size. c) Diversity (left) and novelty (right) of the generated sequences. Higher diversity and novelty is desirable provided that the sequences also have high property value. Diversity is computed as the averaged pairwise hamming distance among the generated sequences. “Labeled” diversity is computed on 500 randomly sampled sequences from the labeled dataset for each protein. For each sequence, novelty is computed as its minimum hamming distance with sequences in the labeled dataset. The novelty of FT generated sequences is very low novelty and is not readily discernible. d) Distributions of oracle predicted property value for the generated sequences. “Labeled” consists of all sequences in the labeled dataset.
Multi-property, Pareto-extrapolative experiment: guiding ESM C for on- and off-target binding of PbrR
Figure 4: Multi-property guidance with ProteinGuide on PbrR to improve lead (Pb) binding and reduce off-target zinc (Zn) binding. a) Schematic diagrams of the data splits used for training and evaluating models for multi-property guidance on PbrR. b) Pareto-extrapolation task: scatterplot of oracle-predicted log-fold change for model generations for ESM C (Unguided), FT, and ProteinGuide. Experimental values of the initial library of variants from [116] are also shown in gray. The target region requires a Pb log-FC of 2 compared to the wild-type and a Zn log-FC of -0.6 compared to wild-type. c) Left panel shows the proportion of sequences that were in the target region for each of ESM C (Unguided), FT, and ProteinGuide. Right panel shows the average Hamming distance between sequences generated by each model, as a measure of diversity. d) Controllable generation task: the top left panel shows outliers and the region corresponding to the remaining four plots (in yellow shading). Top middle panel shows each of three target regions for the Pareto sampling task as quadrants delineated by colored dashed lines. Each target region has it’s bottom left vertex anchored on the approximate Pareto frontier. The remaining three plots (top right and bottom row) show sequences generated by guidance for each target region using the classifier thresholds in the subpanel titles. Subpanel titles also show success rate in parentheses.
Finally in our in silico series of experiments, we tackled a design problem comprising optimizing for two desired properties in tension with one another. This setting is of great interest to protein engineers who may want to, say, increase the activity of an enzyme while simultaneously keeping its stability above some threshold. Moreover, practitioners may want to turn a “knob” to trace out the optimality (i.e., Pareto) frontier in order to characterize it, or to optimize simultaneously for even more than two properties.
Our multi-property optimization task uses data from a recent design campaign [117] that optimized the PbrR protein for use in a cell-free biosensor to detect lead (Pb) contamination in water. PbrR is a transcription factor that forms a dimer when bound to metal ions—typically Pb (II)—allowing it to bind DNA and induce the transcription of a downstream reporter gene. The key challenge is that increasing Pb affinity typically also increases off-target binding to zinc (Zn). We therefore aimed to generate PbrR variants to simultaneously increase Pb binding while reducing off-target Zn (II) binding (Fig. 4a). That is, the goal was to push the Pareto frontier of the experimentally collected data outward past its limits of optimality (Figure 4d).
To simulate the early stage of a design campaign, we only used experimental data from the first library (of five) collected by Wang et al. [116]. This first library comprised of 1,098 variants arising from an alanine scan of the protein; a site saturation mutagenesis (SSM) at 49 positions; and a modest number of higher order mutations (up to degree 5) sampled from seven positions. We only designed positions that had SSM data available, leaving the rest at their wild-type amino acids. Binding affinities are reported as log fold change (log-FC) of fluorescence with respect to the wild-type protein in the presence of its metal-ion cofactor.
To use ProteinGuide we trained two separate predictive models for each of Pb and Zn binding affinities, using all single site and pairwise terms in a linear additive model. We did so to explicitly mimic the common scenario in protein design where initial screens do not capture higher order epistatic interactions and predictive models cannot be expected to learn these interactions from the data. We used ESM C [31] as the pre-trained model. Fine-tuning was done on variants lying on the experimentally observed Pareto frontier for on-/off-target activity (Supplementary Figure S5b, center panel).
We evaluated design for two different tasks. In the first task—Pareto-extrapolation—we aimed to generate sequences in a target region that lay outside the (approximate) Pareto frontier of the experimental data. Specifically, we sought to produce sequences with Pb log-FC greater than $2.0$ and Zn log-FC less than 0.6-0.6 (Fig. 4b, yellow rectangle). In the second task—controllable generation—we selected three reference points along the Pareto frontier representing different Pb-/Zn-binding tradeoffs (Fig. 4d, top-left panel, points on the blue line). For each reference point, the target region corresponded to the part of function space with strictly better Pb and Zn binding characteristics than the reference point (see quadrants delineated by colored dashed lines in Fig. 4d, top-left panel).
For the Pareto-extrapolation task, 75% of the sequences generated using ProteinGuide fell into the design target region (Fig. 4b), without losing any diversity compared to the unguided model (Fig. 4c). The unguided model’s samples had high variance, partially covering the target region (Supplementary Figure S4a). Accordingly, post hoc filtering the unguided sequences (top 100 sequences out of 1,000) yielded only a 16% success rate. The fine-tuned model remained very close to the observed Pareto-optimal frontier in the dataset (i.e., its training data, see Supplementary Figure S5 center panel) and failed to extrapolate (Fig. 4b). Moreover, sequences generated from the fine-tuned model displayed poor diversity (Fig. 4c). In fact, the finetuned model actually predicts lower log-likelihoods for sequences in the target region than the original unconditional model (Supplementary Figure S4f).
Our next experiment—the controllable generation task—showcased the ability of ProteinGuide to conditionally sample sequences according to a diverse range of property target values without retraining generative or predictive models. This is a key advantage of on-the-fly methods as part of design-build-test-learn (DBTL) cycles. As engineered proteins are evaluated in different contexts, the original design criteria may change. With ProteinGuide, new sequences can then be sampled by simply inputting the new property constraints. For methods like fine-tuning, new datasets must be collected and new models must be trained and validated before sequences can be sampled.
In this experiment, we used ProteinGuide to sample $k=50$ sequences for each of three target regions (Fig. 4d, top-right, bottom-left, and bottom-right subpanels). Each target region is anchored on a point from the Pareto frontiers. These anchor points represent different possible lower-bounds on the desired on-/off-target binding affinities for the generated sequences that could be specified by a user during a DBTL cycle. Overall, ProteinGuide had an average success rate of 85% across these three regions showing that it could readily handle a diverse set of user-specified design objectives.
In vivo experiments: engineering a base editor with enhanced activity
[…]